# Circumstellar Interaction in the Ultraviolet Spectra of SN 2023ixf 14–66 Days After Explosion

### v1

@@ 4 Comparison with Literature Models
Although by day 14 the narrow emission lines typically associated with strong CSM interaction have faded, it is possible that some energy from the interaction of the supernova ejecta with CSM is still contributing to the observed supernova spectrum. When supernova ejecta encounter CSM, a cool dense shell is formed between the forward and reverse shocks. The cool dense shell absorbs and reemits some of the shock power from CSM interaction. Dessart & Hillier (2022) published 1D (spherically symmetric) models of the interaction of supernova ejecta and cool dense shell with a low-density wind ($\rm\dot{M}<10^{-3}$ $\mathrm{M_{\odot}~{}yr^{-1}}$). Unlike those of Dessart et al. (2017), which model interaction in a spherically symmetric dense, confined CSM, these models ignore the hydrodynamics of the CSM interaction in exchange for more accurate radiative transfer. While Dessart & Hillier (2022) compared their models to optical observations, they show that the signatures of this interaction are strongest in the UV, and our spectral series allows us to test these model predictions.
@@
Briefly, these models start with a 15 $\mathrm{M_{\odot}}$ progenitor of solar metallicity evolved with Modules for Experiments in Stellar Astrophysics (MESA Paxton et al., 2011, 2013) which has been exploded with the radiation-hydrodynamics code V1D (Livne, 1993; Dessart et al., 2010) and evolved to 10 days post explosion. At day 10, these observations are mapped to the non-local thermodynamic equilibrium radiation transport code CMFGEN (Hillier & Miller, 1998; Hillier & Dessart, 2012; Dessart et al., 2013; Hillier & Dessart, 2019) assuming homologous expansion after this point. A cool dense shell is simulated by placing 0.1 $\mathrm{M_{\odot}}$ of material at 11,700 $\mathrm{km~{}s^{-1}}$ into the density profile. This velocity was chosen as the shock and cool dense shell are not expected to slow down much in the presence of a low-density CSM over a time scale of weeks (Dessart et al., 2017). The power from CSM-ejecta interaction is then deposited at a constant rate into this region, mimicking the conversion of kinetic energy from the ejecta into radiative energy. Implicit in this implementation is the assumption that the cool dense shell is fully formed during the first few days after explosion and does not grow in mass.
@@
With this setup, they produce models with no power deposited and continuous depositions of $1\times 10^{40}$, $1\times 10^{41}$, $5\times 10^{41}$, $1\times 10^{42}$, $2.5\times 10^{42}$, $5\times 10^{42}$, and $1\times 10^{43}$ $\mathrm{erg~{}s^{-1}}$ from day 15 through at least day 120 (some models are run longer). We emphasize that these models were not created for this specific supernova (or any supernova with early narrow emission lines). One difference between the model setup and SN 2023ixf is the presence of confined dense CSM which produces the flash features observed in the early spectra. Dessart et al. (2017) find that in the presence of dense CSM, the maximum ejecta velocity decreased from 11,500 $\mathrm{km~{}s^{-1}}$ in the low-density scenario to 7200 $\mathrm{km~{}s^{-1}}$ in the high-density scenario (see also 2023Dessart). We thus expect the velocity of the outer ejecta of SN 2023ixf to be slower than the models of Dessart & Hillier (2022). Another consequence of the dense CSM is that the initial conditions of the interaction region of the model may not be representative of the physical conditions of the outer ejecta of SN 2023ixf, leading to a larger mismatch between the observations and the models in the earliest epoch than at later times.
@@
We visually compare each model at a given epoch to our observed flux, starting from the day 14.25 spectrum, and identify the model that best matches the observed flux. The spectra of SN 2023ixf and best-matched models are shown in Figure 2 and the model properties are listed in Table 2. As expected, the model velocity is faster than the observed ejecta velocity leading to blueshifted and broader features. We find that the observations of SN 2023ixf indicate a progressive drop in interaction power over time, which translates into a drop in progenitor mass loss with increasing distance.
@@
Table 2: Characteristics of the CSM around the progenitor of SN 2023ixf assuming $v_{shock}=10000$ $\mathrm{km~{}s^{-1}}$ and $v_{wind}=55$ $\mathrm{km~{}s^{-1}}$.
@@
(yr) ($\mathrm{erg~{}s^{-1}}$) ($\mathrm{g~{}cm^{-3}}$) ($\mathrm{M_{\odot}~{}yr^{-1}}$) (cm) (yr)
@@
Note. — The last two rows bracket the true shock luminosity as the Mg II $\lambda\lambda$2796, 2802 emission lies between $\mathrm{L_{sh}=10^{40}}$ and $\mathrm{L_{sh}=10^{41}}$ $\mathrm{erg~{}s^{-1}}$.
@@
Considering that these models were not produced for this specific supernova, the overall flux level is well-matched and many of the features are reproduced, although at varying strengths. The first two epochs of models are too faint in the NUV and too bright in the optical, a discrepancy which improves with time, possibly indicating that the temperature is too low in the emitting region in the early models. This is corroborated by the lack of He I $\lambda$5875 in the models, whose formation requires higher temperatures (Dessart & Hillier, 2005)
@@
The UV is dominated by a forest of iron-group-element absorption lines which blend together, making it challenging to associate features with the emission or P Cygni profiles of individual species (Bostroem et al., 2023c; Dessart & Hillier, 2005). To facilitate the identification of regions with contributions from different elements, we post-process the CMFGEN models to calculate the spectra at day 14 and day 19, omitting the bound-bound transitions of individual species. The full spectrum is then divided by these spectra to obtain the relative flux contribution of each species. We perform this calculation for H, He, C, N, O, Ne, Na, Mg, Al, Si, S, K, Ca, Sc, Ti, Cr, Fe, Co, and Ni. Comparing the omitted spectrum to the full spectrum for each species, regions were identified as at least partially resulting from an individual species. These are marked in the top panel of each column of Figure 4, which shows the full model spectrum compared to the observed spectrum at the same epoch. Elements that have absorption features with a depth of $>2$% of the total flux are shown in the bottom panel of Figure 4. In these panels the strong influence of the iron-group elements (specifically Fe and Ni on day 14 and Fe, Ni, and Cr on day 19) on the overall spectrum is clear. In particular, we point out that the apparent developing “emission line” at 1924 Å is in fact a window of lower absorption between two prominent iron absorption regions, similar to the “emission line” at 2970 Å identified in the NUV spectra of SN 2022acko on day $\sim 20$ (Bostroem et al., 2023c). Additionally, we show the continuum emission if no absorption is present in the top panel of each column of Figure 4, demonstrating how strongly the spectrum deviates from blackbody emission.
@@
Figure 4: Top: The UV spectra of SN 2023ixf on day 14.25 (left) and 19.25 (right) compared to radiative-transfer models. Absorption lines from individual species are marked below the spectra. In the day 19.25 spectrum, the asterisk on Fe at 1924 Å indicates that this is a window of minimal iron absorption rather than an emission line due to any atomic transition. The individual features are reproduced quite well, albeit at higher velocity in the models. Bottom: The decomposition of the model spectrum at each epoch into species that contribute to the spectrum. Although the individual spectra are offset, the scales of the spectra are consistent, showing the relative intensity of each species. The dominance of iron-group elements (and in particular iron) in the spectra at both epochs is clear.
@@
Another emission line of particular note is Mg II in the day 66 spectrum, which is compared to the CMFGEN models in Figure 5. To account for the higher velocity of the models, we scale the models by the ratio of the blue edge of the feature in the observation to the blue edge in the model which preserves the morphology of the line. The Mg II feature is only present in models with CSM interaction after day $\sim 50$ and rapidly increases in strength for greater interaction power. Our observation at day 66 falls between the models with a shock power of $L_{\rm sh}=1\times 10^{40}$ $\mathrm{erg~{}s^{-1}}$ and $L_{\rm sh}=1\times 10^{41}$ $\mathrm{erg~{}s^{-1}}$ and looks clearly distinct from models with no interaction.
@@
In the left panel of Figure 3, we compare the Mg II and $\mathrm{H\alpha}$ profiles of our observation and the $L_{\rm sh}=1\times 10^{40}$ $\mathrm{erg~{}s^{-1}}$ model (without any shift in wavelength) at day 66. We find that both profiles are similar, although the model extends to bluer wavelengths due to the lack of dense CSM in the early supernova evolution which caused the cool dense shell to form at lower velocities in SN 2023ixf. While the models have a large narrow absorption feature on the blue side of $\mathrm{H\alpha}$ from the blue side of the cool dense shell, no such prominent feature is present in the observed $\mathrm{H\alpha}$ profile. There is a hint of an absorption feature around 9200 $\mathrm{km~{}s^{-1}}$, although this is near the noise threshold of the spectrum and thus we cannot confidently associate it with this feature. On the red edge of $\mathrm{H\alpha}$, the models show a subtle red shelf. With the S/N of our data we cannot verify whether this feature is present.
@@
Figure 5: The Mg II line profiles of SN 2023ixf at day 66 in velocity space (black) compared to models with no power from CSM interaction (cyan) and the two lowest levels of CSM interaction modeled ($L_{\rm sh}=1\times 10^{40}$ $\mathrm{erg~{}s^{-1}}$; blue, $L_{\rm sh}=1\times 10^{41}$ $\mathrm{erg~{}s^{-1}}$; magenta). This feature is present only in those models with CSM interaction, indicating that SN 2023ixf is interacting with CSM on day 66. The velocity scale of the models has been scaled to match the observations.
@@ 5 Comparison with Other Supernovae
We focus our comparison of SN 2023ixf to other supernovae on the small number of Type II supernovae with UV observations, defining a sample of objects with a range of CSM properties and supernova types (see Table 3). Although many of these objects continued to exhibit signs of CSM interaction later in their evolution (e.g., Pooley et al., 2002; Mauerhan & Smith, 2012; Smith et al., 2017), we limit the discussion to their plateau-phase evolution which corresponds to our observations of SN 2023ixf.
@@
SN 1993J was a Type IIb supernova, with the hydrogen lines fading and helium lines increasing in strength soon after explosion (Filippenko et al., 1993). Many lines of evidence indicated interaction with CSM including narrow emission lines (Benetti et al., 1994) and X-ray (Zimmermann et al., 1994) and radio (Pooley & Green, 1993) detections. SN 1979C and SN 1998S are both Type II supernovae with steeply declining light curves and above-average photospheric phase luminosities. They also showed narrow emission lines for weeks after explosion; however, these did eventually disappear (Panagia et al., 1980; Leonard et al., 2000). SN 1979C was also detected at X-ray and radio wavelengths indicating interaction with CSM beyond the nominal RSG wind (Panagia et al., 1980). SN 1980K had a decline rate similar to that of SN 1979C, although it was about a magnitude fainter in the $V$ band. While it did not exhibit narrow emission lines, it was detected in X-rays (Canizares et al., 1982) and radio (Weiler et al., 1992). SN 1987A was a peculiar Type II supernova, originating from a compact, blue supergiant progenitor. No CSM interaction was initially detected, however, months after explosion narrow emission lines appeared as high-energy photons ionized a ring of CSM. SN 1999em, SN 2021yja, and SN 2022wsp are considered fairly normal Type II supernovae (Leonard et al., 2000; Baron et al., 2000; Gal-Yam et al., 2008; Hosseinzadeh et al., 2022; Vasylyev et al., 2022, 2023b), although Hosseinzadeh et al. (2022) note that SN 2021yja was exceptionally blue at early times and Vasylyev et al. (2023a) find that the absorption of $\mathrm{H\alpha}$ and $\mathrm{H\beta}$ in SN 2022wsp was suppressed in the P Cygni profiles which they ascribe to CSM interaction. Finally, SN 2005cs and SN 2022acko were low-luminosity supernovae (Brown et al., 2007; Bostroem et al., 2023c). The only possible sign of CSM interaction in these supernovae is a “ledge” feature in the early spectra which has been attributed to CSM, although other possibilities exist (see Pearson et al., 2023, for a detailed discussion). SN 2021yja also showed the ‘ledge’ feature. Details on the explosion epochs, extinction values, and spectra considered in this analysis are given in Table 3. We note that UV data also exist for SN 2010jl, a superluminous Type IIn supernova. Throughout its evolution, the UV spectrum is dominated by the cool dense shell with narrow spectral features and enhanced flux, in contrast to the other supernovae in this sample where the spectra are dominated by supernova photosphere. We therefore exclude it from our analysis.
@@
In Figure 6 and Figure 7 we show a comparison of the spectra of SN 2023ixf to this sample of UV spectra at day $\sim 10$ and day $\sim 20$, respectively. For this comparison, we exclude SN 1979C and SN 1998S, as their UV flux is significantly higher than that of SN 2023ixf and all models considered in this paper. At day 10, SN 2023ixf is similar in brightness to SN 2021yja. Although it has a lower flux, SN 2022wsp is spectroscopically similar. The Mg II $\lambda\lambda$2796, 2802 doublet is present in absorption in SN 1980K, SN 2021yja, SN 2022wsp, and SN 2022acko, although it is blueshifted to higher velocities in SN 2021yja and SN 2022acko. Additionally, blueward of Mg II, the Fe absorption feature identified by Vasylyev et al. (2023b) in SN 2022wsp is clearly present in SN 2023ixf and SN 1980K. Although not a pronounced absorption trough, it is also possibly present in SN 2021yja and SN 2022acko and blended with Mg II in SN 1999em. Like SN 2023ixf, the UV spectra of SN 1980K, SN 2021yja, and SN 2022wsp all show an absorption complex of Fe and Ni between 1750 Å and 2050 Å. The Fe absorption complex between 2050 Å and 2550 Å, which is notably shallower in SN 2023ixf and SN 1980K than models predict, is fairly well reproduced in SN 2021yja, although all of the observed profiles have less prominent individual troughs than the model.
@@
Figure 6: Top: A comparison of the UV spectrum of SN 2023ixf at day 9 with that of other Type II supernovae at a similar epoch (phase for each supernova given in the panel legend): SN 1980K, SN 1987A, SN 1999em, SN 2005cs, SN 2021yja, SN 2022wsp, and SN 2022acko. At NUV wavelengths, SN 2023ixf is among the brightest in the sample. Bottom: The day 9 (black) and 14 (gray) spectra of SN 2023ixf compared with CMFGEN models having varying degrees of shock power at day 14. Although the majority of UV observations have phases around day 10, day 14 is the first epoch modeled by Dessart & Hillier (2022). We show both SN 2023ixf spectra to demonstrate possible evolution from the spectra in the top panel and the models in the bottom. The day 14 spectrum best matches the model with $L_{\rm sh}=5\times 10^{42}$ $\mathrm{erg~{}s^{-1}}$ below $\sim 3000$ but is brighter than the model at shorter wavelengths. Elements are marked in both panels in black at $-7000$ $\mathrm{km~{}s^{-1}}$ to match the SN 2023ixf absorption.
@@
Figure 7: Top: A comparison of the spectrum of SN 2023ixf on day 19 with that of other Type II supernovae at a comparable epoch (phase for each spectrum given in the panel legend): SN 1980K, SN 1987A, SN 1993J, SN 2021yja, SN 2022wsp, and SN 2022acko. SN 2023ixf remains UV bright, especially when compared with SN 2021yja which has faded to a flux similar to that of SN 2022wsp. The explosion epoch of SN 1999em is uncertain and thus we show the spectrum here, too, to demonstrate its similarity to the other spectra at these epochs. Bottom: The UV spectrum of SN 2023ixf at 19 days (black) compared with the CMFGEN models having varying degrees of shock power at day 20 (colors). Although at optical wavelengths the spectrum best matches the model with $L_{\rm sh}=2.5\times 10^{42}$, below $\sim 2750$ the spectrum best matches the model with $L_{\rm sh}=5\times 10^{42}$. Elements are marked in both panels in black at $-7000$ $\mathrm{km~{}s^{-1}}$ to match the SN 2023ixf absorption. The asterisk denotes the feature caused by a window between two strong iron absorption troughs.
@@
One of the most distinctive features of CSM interaction is excess UV flux. At day 10, we see that SN 1980K has a higher flux than SN 2023ixf, SN 2021yja has a very similar flux while SN 2022wsp is a bit lower, SN 2022acko is significantly lower, and the remaining supernovae have virtually no flux at day 10. We note, however, that both excess UV flux and the slope of the UV SED are highly dependent on the assumed distance, extinction, and explosion epoch of the supernovae. In particular, the redshift-independent distance of NGC 4269 ranges from 4 to 7.8 Mpc (see Van Dyk et al., 2019, for a thorough discussion), significantly affecting the luminosity of SN 1980K. While the distances and thus overall luminosities are uncertain, individual features can indicate different levels of CSM interaction. For example, the development of an Fe absorption complex redward of 3000 Å is not present in SN 1980K, SN 2021yja, and SN 2023ixf. This feature only appears in models with lower levels of CSM interaction. This line does appear to be developing in SN 2022wsp, consistent with the interpretation that the lower flux level is indicative of less CSM interaction. Comparing with the models, this suggests that SN 2021yja and SN 2022wsp have close to $1\times 10^{43}$ $\mathrm{erg~{}s^{-1}}$ interaction power added to the ejecta luminosity at day 10.

### v2

@@ 4 Comparison with Literature Models
Although by day 14 the narrow emission lines typically associated with strong CSM interaction have faded, it is possible that some energy from the interaction of the supernova ejecta with CSM is still contributing to the observed supernova spectrum. When supernova ejecta encounter CSM, a cool dense shell is formed between the forward and reverse shocks. The cool dense shell absorbs and reemits some of the shock power from CSM interaction. Dessart & Hillier (2022) published 1D (spherically symmetric) models of the interaction of supernova ejecta and cool dense shell with a low-density wind ($\rm\dot{M}<10^{-3}$ $M\mathrm{{}_{\odot}~{}yr^{-1}}$). Unlike those of Dessart et al. (2017), which model interaction in a spherically symmetric dense, confined CSM, these models ignore the hydrodynamics of the CSM interaction in exchange for more accurate radiative transfer. While Dessart & Hillier (2022) compared their models to optical observations, they show that the signatures of this interaction are strongest in the UV, and our spectral series allows us to test these model predictions.
@@
Briefly, these models start with a 15 $M_{\mathrm{\odot}}$ progenitor of solar metallicity evolved with Modules for Experiments in Stellar Astrophysics (Paxton et al., 2011, 2013) which has been exploded with the radiation-hydrodynamics code V1D (Livne, 1993; Dessart et al., 2010) and evolved to 10 days post explosion. On day 10, these observations are mapped to the nonlocal thermodynamic equilibrium radiation transport code CMFGEN (Hillier & Miller, 1998; Hillier & Dessart, 2012; Dessart et al., 2013; Hillier & Dessart, 2019) assuming homologous expansion after this point. A cool dense shell is simulated by placing 0.1 $M_{\mathrm{\odot}}$ of material at 11,700 $\mathrm{km~{}s^{-1}}$ into the density profile. This velocity was chosen as the shock and cool dense shell are not expected to slow down much in the presence of a low-density CSM over a time scale of weeks (Dessart et al., 2017). The power from the CSM-ejecta interaction is then deposited at a constant rate into this region, mimicking the conversion of kinetic energy from the ejecta into radiative energy. Implicit in this implementation is the assumption that the cool dense shell is fully formed during the first few days after the explosion and does not grow in mass.
@@
With this setup, they produce models with no power deposited and continuous depositions of $1\times 10^{40}$, $1\times 10^{41}$, $5\times 10^{41}$, $1\times 10^{42}$, $2.5\times 10^{42}$, $5\times 10^{42}$, and $1\times 10^{43}$ $\mathrm{erg~{}s^{-1}}$ from day 15 through at least day 120 (some models are run longer). We emphasize that these models were not created for this specific supernova (or any supernova with early narrow emission lines). One difference between the model setup and SN 2023ixf is the presence of confined dense CSM, which produces the flash features observed in the early spectra. Dessart et al. (2017) find that in the presence of dense CSM, the maximum ejecta velocity decreased from 11,500 $\mathrm{km~{}s^{-1}}$ in the low-density scenario to 7200 $\mathrm{km~{}s^{-1}}$ in the high-density scenario (see also Dessart & Jacobson-Galán, 2023). We thus expect the velocity of the outer ejecta of SN 2023ixf to be slower than the models of Dessart & Hillier (2022). Another consequence of the dense CSM is that the initial conditions of the interaction region of the model may not be representative of the physical conditions of the outer ejecta of SN 2023ixf, leading to a larger mismatch between the observations and the models in the earliest epoch than at later times.
@@
We visually compare each model at a given epoch to our observed flux, starting from the day 14.25 spectrum, and identify the model that best matches the observed flux. Although each of the models represents a steady-state wind, the cool, dense shell is optically thin to continuum radiation and radiates efficiently. For this reason, it will react promptly to a change in shock power. Thus, we expect that a decreasing mass-loss rate with radius would effectively have the same result as taking a different steady-state model for different epochs. The spectra of SN 2023ixf and best-matched models are shown in Figure 2 and the model properties are listed in Table 2. As expected, the model velocity is faster than the observed ejecta velocity leading to blueshifted and broader features. We find that the observations of SN 2023ixf indicate a progressive drop in interaction power over time, which translates into a drop in progenitor mass loss with increasing distance.
@@
Table 2: Characteristics of the CSM around the progenitor of SN 2023ixf assuming $v_{\mathrm{shock}}=10000$ $\mathrm{km~{}s^{-1}}$ and $v_{\mathrm{wind}}=55$ $\mathrm{km~{}s^{-1}}$.
@@
(yr) ($\mathrm{erg~{}s^{-1}}$) ($\mathrm{g~{}cm^{-3}}$) ($M\mathrm{{}_{\odot}~{}yr^{-1}}$) (cm) (yr)
@@
Note. — The last two rows bracket the true shock luminosity as the Mg II $\lambda\lambda$2796, 2802 emission lies between $L\mathrm{{}_{sh}=10^{40}}$ and $L\mathrm{{}_{sh}=10^{41}}$ $\mathrm{erg~{}s^{-1}}$.
@@
Considering that these models were not produced for this specific supernova, it is impressive that no scaling beyond correcting for distance is required to match the observations and many of the features are reproduced, although at varying strengths. The first two epochs of models are too faint in the NUV and too bright in the optical, a discrepancy that improves with time. Nevertheless, we consider the best-matched model to be indicative of the CSM interaction level as the selected model most closely matches the observed flux: the model with more CSM interaction is too bright at all wavelengths and the model with less CSM interaction is even more discrepant for the majority of the spectrum. This consideration is further validated by the custom modeling in Section 6. One possible reason for the difference between the models and observations is that the temperature is too low in the emitting region in the early models. This is corroborated by the lack of He I $\lambda$5875 in the models, whose formation requires higher temperatures (Dessart & Hillier, 2005)
@@
The UV is dominated by a forest of iron-group-element absorption lines that blend together, making it challenging to associate features with the emission or P Cygni profiles of individual species (Bostroem et al., 2023c; Dessart & Hillier, 2005). To facilitate the identification of regions with contributions from different elements, we post-process the CMFGEN models to calculate the spectra on day 14 and day 19, omitting the bound-bound transitions of individual species. The full spectrum is then divided by these spectra to obtain the relative flux contribution of each species. We perform this calculation for H, He, C, N, O, Ne, Na, Mg, Al, Si, S, K, Ca, Sc, Ti, Cr, Fe, Co, and Ni. Comparing the omitted spectrum to the full spectrum for each species, regions were identified as at least partially resulting from an individual species. These are marked in the top panel of each column of Figure 4, which shows the full model spectrum compared to the observed spectrum at the same epoch. Elements that have absorption features with a depth of $>2$% of the total flux are shown in the bottom panel of Figure 4. In these panels, the strong influence of the iron-group elements (specifically Fe and Ni on day 14 and Fe, Ni, and Cr on day 19) on the overall spectrum is clear. In particular, we point out that the apparent developing “emission line” at 1924 Å is in fact a window of lower absorption between two prominent iron absorption regions, similar to the “emission line” at 2970 Å identified in the NUV spectra of SN 2022acko on day $\sim 20$ (Bostroem et al., 2023c). Additionally, we show the continuum emission if no absorption is present in the top panel of each column of Figure 4, demonstrating how strongly the spectrum deviates from blackbody emission.
@@
Figure 4: Top: UV spectra of SN 2023ixf on day 14.25 (left) and 19.25 (right) compared to radiative-transfer models. Absorption lines from individual species are marked below the spectra. In the day 19.25 spectrum, the asterisk on Fe at 1924 Å indicates that this is a window of minimal iron absorption rather than an emission line due to any atomic transition. The individual features are reproduced quite well, albeit at higher velocities in the models. Bottom: decomposition of the model spectrum at each epoch into species that contribute to the spectrum. Although the individual spectra are offset, the scales of the spectra are consistent, showing the relative intensity of each species. The dominance of iron-group elements (and in particular iron) in the spectra at both epochs is clear.
@@
Another emission line of particular note is Mg II in the day 66 spectrum, which is compared to the CMFGEN models in Figure 5. To account for the higher velocity of the models, we scale the models by the ratio of the blue edge of the feature in the observation to the blue edge in the model which preserves the morphology of the line. The Mg II feature is only present in models with CSM interaction after day $\sim 50$ and rapidly increases in strength for greater interaction power. Our observation on day 66 falls between the models with a shock power of $L_{\rm sh}=1\times 10^{40}$ $\mathrm{erg~{}s^{-1}}$ and $L_{\rm sh}=1\times 10^{41}$ $\mathrm{erg~{}s^{-1}}$ and looks clearly distinct from models with no interaction.
@@
In the left panel of Figure 3, we compare the Mg II and $\mathrm{H\alpha}$ profiles of our observation and the $L_{\rm sh}=1\times 10^{40}$ $\mathrm{erg~{}s^{-1}}$ model (without any shift in wavelength) on day 66. We find that both profiles are similar, although the model extends to bluer wavelengths due to the lack of dense CSM in the early supernova evolution, which caused the cool dense shell to form at lower velocities in SN 2023ixf. While the models have a large narrow absorption feature on the blue side of $\mathrm{H\alpha}$ from the blue side of the cool dense shell, no such prominent feature is present in the observed $\mathrm{H\alpha}$ profile. There is a hint of an absorption feature around 9200 $\mathrm{km~{}s^{-1}}$, although this is near the noise threshold of the spectrum and thus we cannot confidently associate it with this feature. On the red edge of $\mathrm{H\alpha}$, the models show a subtle red shelf. With the S/N of our data, we cannot verify whether this feature is present.
@@
Figure 5: Mg II line profiles of SN 2023ixf on day 66 in velocity space (black) compared to models with no power from CSM interaction (cyan) and the two lowest levels of CSM interaction modeled ($L_{\rm sh}=1\times 10^{40}$ $\mathrm{erg~{}s^{-1}}$; blue, $L_{\rm sh}=1\times 10^{41}$ $\mathrm{erg~{}s^{-1}}$; magenta). This feature is present only in those models with CSM interaction, indicating that SN 2023ixf is interacting with CSM on day 66. The velocity scale of the models has been scaled to match the observations.
@@ 5 Comparison with Other Supernovae
We focus our comparison of SN 2023ixf to other supernovae on the small number of Type II supernovae with UV observations, defining a sample of objects with a range of CSM properties and supernova types (see Table 3). Although many of these objects continued to exhibit signs of CSM interaction later in their evolution (e.g., Pooley et al., 2002; Mauerhan & Smith, 2012; Smith et al., 2017), we limit the discussion to their plateau-phase evolution, which corresponds to our observations of SN 2023ixf.
@@
SN 1993J was a Type IIb supernova, with the hydrogen lines fading and helium lines increasing in strength soon after explosion (Filippenko et al., 1993). Many lines of evidence indicated interaction with CSM including narrow emission lines (Benetti et al., 1994) and X-ray (Zimmermann et al., 1994) and radio (Pooley & Green, 1993) detections. SN 1979C and SN 1998S are both Type II supernovae with steeply declining light curves and above-average photospheric phase luminosities. They also showed narrow emission lines for weeks after explosion; however, these did eventually disappear (Panagia et al., 1980; Leonard et al., 2000). SN 1979C was also detected at X-ray and radio wavelengths indicating interaction with CSM beyond the nominal RSG wind (Panagia et al., 1980). SN 1980K had a decline rate similar to that of SN 1979C, although it was about a magnitude fainter in the $V$ band. While it did not exhibit narrow emission lines, it was detected in X-rays (Canizares et al., 1982) and radio (Weiler et al., 1992). SN 1987A was a peculiar Type II supernova, originating from a compact, blue supergiant progenitor. No CSM interaction was initially detected; however, months after explosion narrow emission lines appeared as high-energy photons ionized a ring of CSM. SN 1999em, SN 2021yja, and SN 2022wsp are considered fairly normal Type II supernovae (Leonard et al., 2000; Baron et al., 2000; Gal-Yam et al., 2008; Hosseinzadeh et al., 2022; Vasylyev et al., 2022, 2023a), although Hosseinzadeh et al. (2022) note that SN 2021yja was exceptionally blue at early times and Vasylyev et al. (2023b) find that the absorption of $\mathrm{H\alpha}$ and $\mathrm{H\beta}$ in SN 2022wsp was suppressed in the P Cygni profiles which they ascribe to CSM interaction. Finally, SN 2005cs and SN 2022acko were low-luminosity supernovae (Brown et al., 2007; Bostroem et al., 2023c). The only possible sign of CSM interaction in these supernovae is a “ledge” feature in the early spectra, which has been attributed to CSM, although other possibilities exist (see Pearson et al., 2023, for a detailed discussion). SN 2021yja also showed the ‘ledge’ feature. Details on the explosion epochs, extinction values, and spectra considered in this analysis are given in Table 3. We note that UV data also exist for SN 2010jl, a superluminous Type IIn supernova. Throughout its evolution, the UV spectrum is dominated by the cool dense shell with narrow spectral features and enhanced flux, in contrast to the other supernovae in this sample where the spectra are dominated by the supernova photosphere. We therefore exclude it from our analysis.
@@
In Figure 6 and Figure 7 we show a comparison of the spectra of SN 2023ixf to this sample of UV spectra, scaled to the distance of SN 2023ixf, on day $\sim 10$ and day $\sim 20$, respectively. For this comparison, we exclude SN 1979C and SN 1998S, as their UV flux is significantly higher than that of SN 2023ixf and all models considered in this paper. On day 10, SN 2023ixf is similar in brightness to SN 2021yja. Although it has a lower flux, SN 2022wsp is spectroscopically similar. The Mg II $\lambda\lambda$2796, 2802 doublet is present in absorption in SN 1980K, SN 2021yja, SN 2022wsp, and SN 2022acko, although it is blueshifted to higher velocities in SN 2021yja and SN 2022acko. Additionally, blueward of Mg II, the Fe absorption feature identified by Vasylyev et al. (2023a) in SN 2022wsp is clearly present in SN 2023ixf and SN 1980K. Although not a pronounced absorption trough, it is also possibly present in SN 2021yja and SN 2022acko and blended with Mg II in SN 1999em. Like SN 2023ixf, the UV spectra of SN 1980K, SN 2021yja, and SN 2022wsp all show an absorption complex of Fe and Ni between 1750 Å and 2050 Å. The Fe absorption complex between 2050 Å and 2550 Å, which is notably shallower in SN 2023ixf and SN 1980K than models predict, is fairly well reproduced in SN 2021yja, although all of the observed profiles have less prominent individual troughs than the model.
@@
Figure 6: Top: comparison of the UV spectrum of SN 2023ixf on day 9 with that of other Type II supernovae at a similar epoch (phase for each supernova given in the panel legend; all spectra scaled to the distance of SN 2023ixf): SN 1980K, SN 1987A, SN 1999em, SN 2005cs, SN 2021yja, SN 2022wsp, and SN 2022acko. At NUV wavelengths, SN 2023ixf is among the brightest in the sample. Bottom: day 9 (black) and 14 (gray) spectra of SN 2023ixf compared with CMFGEN models having varying degrees of shock power on day 14. Although the majority of UV observations have phases around day 10, day 14 is the first epoch modeled by Dessart & Hillier (2022). We show both SN 2023ixf spectra to demonstrate possible evolution from the spectra in the top panel and the models in the bottom. The day 14 spectrum best matches the model with $L_{\rm sh}=5\times 10^{42}$ $\mathrm{erg~{}s^{-1}}$ below $\sim 3000$ but is brighter than the model at shorter wavelengths. Elements are marked in both panels in black at $-7000$ $\mathrm{km~{}s^{-1}}$ to match the SN 2023ixf absorption.
@@
Figure 7: Top: comparison of the spectrum of SN 2023ixf on day 19 with that of other Type II supernovae at a comparable epoch (phase for each spectrum given in the panel legend; all spectra scaled to the distance of SN 2023ixf): SN 1980K, SN 1987A, SN 1993J, SN 2021yja, SN 2022wsp, and SN 2022acko. SN 2023ixf remains UV bright, especially when compared with SN 2021yja, which has faded to a flux similar to that of SN 2022wsp. The explosion epoch of SN 1999em is uncertain and thus we show the spectrum here, too, to demonstrate its similarity to the other spectra at these epochs. Bottom: UV spectrum of SN 2023ixf on 19 days (black) compared with the CMFGEN models having varying degrees of shock power on day 20 (colors). Although at optical wavelengths the spectrum best matches the model with $L_{\rm sh}=2.5\times 10^{42}$, below $\sim 2750$ the spectrum best matches the model with $L_{\rm sh}=5\times 10^{42}$. Elements are marked in both panels in black at $-7000$ $\mathrm{km~{}s^{-1}}$ to match the SN 2023ixf absorption. The asterisk denotes the feature caused by a window between two strong iron absorption troughs.
@@
One of the most distinctive features of CSM interaction is excess UV flux. On day 10, we see that SN 1980K has a higher flux than SN 2023ixf, SN 2021yja has a very similar flux while SN 2022wsp is a bit lower, SN 2022acko is significantly lower, and the remaining supernovae have virtually no flux on day 10. We note, however, that both excess UV flux and the slope of the UV SED are highly dependent on the assumed distance, extinction, and explosion epoch of the supernovae. In particular, the redshift-independent distance of NGC 4269 ranges from 4 to 7.8 Mpc (see Van Dyk et al., 2019, for a thorough discussion), significantly affecting the luminosity of SN 1980K. While the distances and thus overall luminosities are uncertain, individual features can indicate different levels of CSM interaction. For example, the development of an Fe absorption complex redward of 3000 Å is not present in SN 1980K, SN 2021yja, and SN 2023ixf. This feature only appears in models with lower levels of CSM interaction. This line does appear to be developing in SN 2022wsp, consistent with the interpretation that the lower flux level is indicative of less CSM interaction. Comparing with the models, this suggests that SN 2021yja and SN 2022wsp have close to $1\times 10^{43}$ $\mathrm{erg~{}s^{-1}}$ interaction power added to the ejecta luminosity on day 10.